# Decentralized Identity Management on Ripple: A Conceptual Framework for High-Speed, Low-Cost Identity Transactions in Attestation-Based Attribute-Based Identity

### v1

@@ 3.1.2 Example Scenarios
    Scenario 2: A certain credit facility is offered by Bank of A to new or old customers with at least 70,000\$70,000\\$ liquidity at a reputable institution. This "Reputation" of the issuer is an important attribute for a verifier in this niche of ABIDMs. Since Ripple’s partners, current and potential, are reputed financial institutions and government authorities, metrics in our IDM for the reputation we deem unnecessary. A customer with more than 40,000\$40,000\\$ at Bank of B and more than 55,000\$55,000\\$ at Bank of C has statements "More than 25,000\$25,000\\$ at B" and "More than 45,000\$45,000\\$ at C." The respective banks issue attestations for these claims. Verifier A can see that the customer exceeds the liquidity threshold without discovering their actual amount of liquid funds at other institutions; the customer has no funds at A since he is a new customer.
@@ 3.1.4 Scenario 2
The banks of B and C have the statements given previously. Both institutions know that these statements are valid. Identical to the above scenario, B initiates the transaction and signs the statement, followed by the customer’s signature, and corresponding transactions are executed with $0.3XRP$ as the transaction volume. The same takes place regarding the attestation from C. The addresses of both these attestation issuance transactions are stored on the customer’s account by the IC, which he shares with the Bank of A. The VC would pull and display these transaction details: statements; attestations; signature verifications; and the identities of B, C, and the customer. By verifying the legitimacy of the attestation, i.e., the double signature verifications, it is confirmed by A that the customer has liquidity more than 70,000\$70,000\\$. Hence, the credit facility can be approved for the customer.

### v2

@@ 1 Introduction
1.1 Revised Contribution in Light of New Decentralized Identifiers (DID) on Ripple
In light of the recent XLS-40[5] amendment on the XRP Ledger, which introduces native support for W3C DIDs, this work still makes contributions by advancing a layered identity-management design rather than merely defining identifiers. While XRPL’s DID method allows on-ledger storage or references of DID documents, it does not inherently address the full lifecycle of verifiable credentials, attribute attestations, revocation mechanisms, or privacy-preserving attribute disclosure. Our conceptual framework proposes an Attestation-Based, Attribute-Based Decentralized Identity Management (ABABDIDM) system that defines issuers, verifiers, and end-users in a high-throughput, low-cost architecture. We show how attribute attestations can be efficiently managed, how credential revocation can be supported, a sidechain for identity storage, a cross-chain interoperability approach, XRPL(and sidechain)-EVM bridging for logic execution, and how trust anchors might be governed, all grounded on XRPL’s performance and consensus structure. This design provides a path toward scalable and usable identity systems that go beyond basic DID resolution.
@@ 3.1.2 Example Scenarios
    Scenario 2: A certain credit facility is offered by Bank of A to new or old customers with at least $70,000\mathdollar$ liquidity at a reputable institution. This "Reputation" of the issuer is an important attribute for a verifier in this niche of ABIDMs. Since Ripple’s partners, current and potential, are reputed financial institutions and government authorities, metrics in our IDM for the reputation we deem unnecessary. A customer with more than $40,000\mathdollar$ at Bank of B and more than $55,000\mathdollar$ at Bank of C has statements "More than $25,000\mathdollar$ at B" and "More than $45,000\mathdollar$ at C." The respective banks issue attestations for these claims. Verifier A can see that the customer exceeds the liquidity threshold without discovering their actual amount of liquid funds at other institutions; the customer has no funds at A since he is a new customer.
@@ 3.1.4 Scenario 2
The banks of B and C have the statements given previously. Both institutions know that these statements are valid. Identical to the above scenario, B initiates the transaction and signs the statement, followed by the customer’s signature, and corresponding transactions are executed with $0.3XRP$ as the transaction volume. The same takes place regarding the attestation from C. The addresses of both these attestation issuance transactions are stored on the customer’s account by the IC, which he shares with the Bank of A. The VC would pull and display these transaction details: statements; attestations; signature verifications; and the identities of B, C, and the customer. By verifying the legitimacy of the attestation, i.e., the double signature verifications, it is confirmed by A that the customer has liquidity more than $70,000\mathdollar$. Hence, the credit facility can be approved for the customer.
@@ 4 Future Directions
4.1 Limitations
While the proposed Attestation-Based, Attribute-Based Decentralized Identity Management (ABABDIDM) framework builds on XRPL’s XLS-40 DID capabilities and outlines a conceptual pathway for cross-chain identity management, several limitations must be acknowledged:
  1. 1.
    Formal Security Analysis Not Included. This work is conceptual and does not provide formal proofs for security against adversarial scenarios, such as colluding issuers, malicious verifiers, or sidechain compromise.
  2. 2.
    Privacy Guarantees are Conceptual. Although the framework considers privacy-preserving attribute disclosure, detailed protocols (e.g., zero-knowledge proofs) are not implemented, and full unlinkability is not formally established.
  3. 3.
    Operational Architecture is Abstract. The IDM sidechain, cross-chain interoperability, and XRPL- and sidechain-EVM bridging are described at a high level; deployment strategies, transaction throughput, and cost evaluations remain future work.
  4. 4.
    Trust and Governance. While the framework outlines trust anchors and governance mechanisms, detailed incentive structures and policies for issuers, verifiers, and end-users are not fully developed.
  5. 5.
    Revocation and Lifecycle Management. The approach to credential revocation and lifecycle management is proposed conceptually; efficient on-chain/off-chain revocation mechanisms require further research.
  6. 6.
    Dependency on XRPL Infrastructure. The framework assumes XLS-40 DID infrastructure and XRPL ledger stability; future changes to XRPL consensus or DID standards may require adaptation of the proposed system.
  7. 7.
    Preliminary Conceptual Stage. The work is meant to communicate the architecture and design principles for a high-throughput, low-cost identity system. Full implementation, testing, and benchmarking are deferred to future research.